Circular No. 8067

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003af Further to IAUC 8066, W. M. Wood-Vasey, G. Aldering, and P. Nugent

report the discovery of an apparent supernova (mag 15.6) on unfiltered NEAT images taken on Feb. 3.56 and 3.57 UT. The object appeared at mag 15.7 on images from Feb. 7. The new object is located at α = 11^{h}10^{m}21^{s}.69, δ = +23^{o}24′50′′.9 (equinox 2000.0), which is 4′′ east and 2′′ south of the center of the apparent host galaxy. There is evidence for a point source at this location on images taken with the same telescope on Jan. 19, but nothing is present in Palomar images from the Digital Sky Survey.

SUPERNOVA 2003ad T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003ad (cf. IAUC 8065), obtained by M. Calkins on Feb. 6.53 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova soon after explosion. The spectrum consists of a very blue continuum with incipient lines of hydrogen. A weak, narrow Hα emission line superposed on the spectrum of the supernova indicates a recession velocity of 10500 km/s for the host galaxy. The supernova expansion velocity, derived from the minimum of the Hβ line, is 10100 km/s.

COMET C/2003 A2 (GLEASON) Additional observations (including prediscovery observations from last

November) and the following osculating hyperbolic orbital elements are taken from MPEC 2003-C47. The “original” barycentric value of 1/a is −0.00042 ± 0.00027 AU^{−1} (1σ mean error).

Epoch = 2003 Dec. 27.0 TT

T = 2004 Jan. 15.62696 TT  ω = 349.^{o}32111
e = 1.0121852  Ω = 154.49031 }2000.0
q = 11.3814634 AU  i = 8.07943

COMET C/2002 X5 (KUDO-FUJIKAWA) Visual m_{1}, coma-diameter, and tail-length estimates: Feb. 7.51 UT,

5.7, 2′, > 0^{o}.5 in p.a. 158^{o} (M. Mattiazzo, Wallaroo, S. Australia, 25×100 binoculars; altitude 3^{o}); 8.51, 6.0, 3′, 2^{o}.0 (A. Pearce, City Beach, W. Australia, 20×80 binoculars).

2003 February 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green